Recent Maintenance – KURK

KURK operational status:


The station operator replaced the KURK-00 Q330 digitizer (replacing Q330 that was experiencing reboots).


IDA engineers visited the station and replaced both seismometers, the strong motion sensor, and upgraded the power system.  New sensors include an STS-6, STS-5, and an Episensor strong motion sensor.


An IDA engineer visited the station to upgrade the power system and install Setra and Paroscientific microbarometers.


The station is up as of day 2017,038 after repairs to a damaged surge suppressor and new fuses were installed.


The station is down as of day 2017,031 due to a power surge causing damages to the power system.


There are telemetry issues with the station since March 2015. An IDA engineer is working with the station operator to resolve these issues. The station operator is posting data from the station every 2 weeks for IDA to retrieve until normal telemetry is restored.


An IDA engineer is currently on-site upgrading the station to a Quanterra Q330HR system. Sensors include STS-1 and STS-2 seismometers, and a strong motion sensor.


The station is back on-line as of day 2010,253 after the station operator replaced the GPS clock and the XCOM board in the DAS.


The station is acquiring data again after the station operator replaced the DAS with the previous DAS used at the station. However, there is a timing problem and the data have no time stamps. IDA engineers are troubleshooting the problem and will contact the station operator.


The station is down as of day 2010,207-09:14 after a thunderstorm. Plans are being made for a station visit to upgrade the station and restore data acquisition.


The station is back on-line after the station operator replaced the DAS. The station operator had replaced the GPS clock on July 28 but the station remained down due to problems with the DAS.


The GPS clock at KURK failed about day 2009,178-15:24. A replacement GPS clock is being prepared to ship to the station.


An IDA engineer visited the station and replaced the East-West and North-South STS-1 sensors, the East-West GS-13 sensor, and the batteries for the UPS. This restored data for the East-West short period channels and improved the data quality for the horizontal broadband channels.


The station is back up after the station operator rebooted the ISP.


The station is down due to problems with the ISP.


The station operator reports (8/3/2004) that the East-West channels from the STS-1 sensor are noisy, after the masses for the North-South and Vertical channels were recentered. The data quality for the EW and NS channels for the STS-1 sensor is poor.


Station back up on 2003.12.05, after an IDA engineer replaced the station’s power supply and XCOM board in the digitizer.


DC-DC converter shipped to station (10/03).


Station operator reports replacing power supply shipped 9/4/2003, however, the station remains down due to problems with the remote power supply.


Station down due to main power supply failure. No data collected since 211-04:41:59 (July 30,2003).


The EW component of the GS-13 sensor began malfunctioning on this day.


Station is now operational after replacement XCOM circuit board was installed. New EPROMs installed to correct DSP problem.


Station is currently down due to lightning damage on July 16th.


Station operating normally. The NS STS1 sensor is noisy.